OpenStudy (anonymous):

if you are selecting a random sample of n=5 marbles and the first 4 marbles are all blue, what is the probability that the fifth marble will be green?

OpenStudy (dumbcow):

i dont know, how many blue and green marbles are there?

OpenStudy (dumbcow):

if you are looking for prob that the sample is 4 blue 1 green and the the prob of picking blue is 1/2 then use binomial distribution P = 5C4*(1/2)^4*(1/2) = 5*(1/32) = 5/32

OpenStudy (anonymous):

It's unclear... You must state how many green marbles there are surely?

OpenStudy (anonymous):

A binomial distribution implies there are only two variables or numbers. Hence bi- (two) and -nomial (relating to numbers). The Binomial Distribution will only hold true if there are green and blue marbles exclusively. It's not specified in the question so the answer is unclear surely?